OpenStudy (anonymous):

what is the second derivative of y with respect to x of 4x^2 + 9y^2 = 36? a.16/9y^2 b.-(9/16y^2) c.9/16y^2 d.-(16/9y^2)

OpenStudy (anonymous):

JAMES i want to put letter E here

OpenStudy (anonymous):

e for error

OpenStudy (anonymous):

start with \[8x+18yy'=0\] solve for y'

OpenStudy (jamesj):

Yes, and then differentiate again.

OpenStudy (anonymous):

oops second derivative if i could learn to read i would not need math...

OpenStudy (anonymous):

what jamesj said this time you will need the quotient rule

OpenStudy (anonymous):

you got this or you need the steps? because i think there will be something to do here

OpenStudy (anonymous):

w8 i put it on the notebook this is the quiz tomorrow.

OpenStudy (anonymous):

is this step \[8x+18yy'=0\] clear?

OpenStudy (anonymous):

how can i translate that word in english w8 w8 = u doble it then theres a y prime before 0 because of y^2

OpenStudy (anonymous):

i am taking the derivative of \[y^2\] with respect to x, using the chain rule and thinking that y is a function of x

OpenStudy (anonymous):

even though we don't have it written as one. that is why it is called "implicit" diff

OpenStudy (anonymous):

for example if \[y^2=\sin^2(x)\] then \[(y^2)'=2\sin(x)\cos(x)\] and if \[y^2=f^2(x)\] then \[(y^2)'=2f(x)f'(x)\] so in general \[(y^2)'=2yy'\]

OpenStudy (anonymous):

this is the last question on my paper explain it more satelite i need it.

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

think of y as \[y=f(x)\]

OpenStudy (anonymous):

then \[y'=f'\] and if you have \[y^2\] then you need the chain rule

OpenStudy (anonymous):

now we solve for y' and that is just algebra \[8x+18yy'=0\] \[y'=-\frac{4x}{9y}\]

OpenStudy (anonymous):

now we need \[y''\] so take the derivative again, this time using the quotient rule

OpenStudy (anonymous):

then y'' = - 8x/18yy'

OpenStudy (anonymous):

again we think of \[y=f(x)\]

OpenStudy (anonymous):

so it is like taking the derivative of \[-\frac{4x}{9f(x)}\]

OpenStudy (anonymous):

the denominator is easy, just square it

OpenStudy (anonymous):

lets pull out the constant first so we have \[-\frac{4}{9}\frac{x}{y}\]

OpenStudy (anonymous):

then take the derivative of \[\frac{x}{y}\] which again is like taking the derivative of \[\frac{x}{f(x)}\] you get \[\frac{f(x)-x\times f'(x)}{f^2(x)}\] or \[\frac{y-xy'}{y^2}\]

OpenStudy (anonymous):

therefore we have our derivative is \[-\frac{4}{9}\frac{y-xy'}{y^2}\] now come the only tricky part

OpenStudy (anonymous):

thats y u get the 9y^2?

OpenStudy (anonymous):

we are not there yet

OpenStudy (anonymous):

you see that \[y'\] in the numerator, we have to get rid of it

OpenStudy (anonymous):

replace the y' in the numerator by \[-\frac{4x}{y}\] because that is what we found it to be earlier then do some algebra so simplify the compound fraction

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

question sir is x = 1 ryt? 4(1)^2

OpenStudy (anonymous):

\[-\frac{4}{9}\times \frac{y-x(\frac{-4x}{9y})}{y^2}\] multiply top and bottom by \[9y\] and get \[-\frac{9y^2+4x^2}{9y^3}\]

OpenStudy (anonymous):

actually you get \[-\frac{4}{9}\frac{9y^2+4x^2}{9y^3}\]

OpenStudy (anonymous):

theres some1 message on me that my answer about this link he said my answer and your answer is wrong http://openstudy.com/study#/updates/4f561b47e4b0862cfd072588 its from another site.

OpenStudy (anonymous):

but then you recall from the orignal question that \[9y^2+4x^2=36\] so replace the numerator by 36 and you are done

OpenStudy (anonymous):

i replace the numerator by 36

OpenStudy (anonymous):

yup, and that should do it!

OpenStudy (anonymous):

can i copy this to my notebook?

OpenStudy (anonymous):

i just need it

OpenStudy (anonymous):

sure why not?

OpenStudy (anonymous):

answer is letter b i do what u do

OpenStudy (anonymous):

ok but i think the denomiator should have y^3 in it, guess i am wrong

OpenStudy (anonymous):

i mean i really think i am right, but maybe i made a mistake somewhere in any case i have to run

OpenStudy (anonymous):

my prof is very tricky where do you think he get all this question? from different books?

OpenStudy (anonymous):

he always give students a xerox of problem set like this =)